OpenStudy (anonymous):

You have to pay the local council a service fee of 25 dollars for the water supply and 80 cents per kilolitre of water used, Dies the amount of water you pay for directly vary with the amount of water used?

OpenStudy (anonymous):

My answer is: Yes, it does as the fee varies according to the amount of water used. eg. f = 25 + 0.8(1) = 25.8 and f = 25 + 08(2) = 26.6